OpenStudy (anonymous):

The probability of a student scoring 75% in class work is 0.64, and the probability of a student scoring 85% is 0.45. Event A: The student scores 75%. Event B: The student scores 85%. The probability of a student scoring 85% in class work, given that they have already scored 75% in class work, is 0.55. The probability of a student scoring 75% in class work, given that they have already scored 85% in class work, is 1. Which statement is true?

OpenStudy (anonymous):

Events A and B are independent because P(A|B) = P(A). Events A and B are independent because P(A|B) = P(B). Events A and B are independent because P(A|B) = P(A) + P(B). Events A and B are not independent because P(A| B) ≠ P(A). Events A and B are not independent because P(A|B) = P(A).these are my choices